U have to Prove that: cos π/7 *cos 2π/7 *cos 3π/7 =1/8

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
A = cos 3π/7. cos 2π/7. cos π/7 
cos 3π/7 = cos 6π/14 = cos ( π - 8π/14 ) = - cos 8π/14 
Then, 
A = - cos 8π/14. cos 4π/14. cos 2π/14 
= - cos x. cos 2x. cos 4x, ... x = 2π/14 
= (-1/ 2 sin x ). ( 2 sin x. cos x ). cos 2x. cos 4x 
= (-1/ 4 sin x ). ( 2 sin 2x. cos 2x ). cos 4x 
= (-1/ 8 sin x ). ( 2 sin 4x. cos 4x ) 
= (-1/8). ( 1 / sin x ). sin 8x 
= (-1/8). ( 1/ sin π/7 ). sin 8π/7 
= (-1/8). ( 1/ sin π/7 ). sin ( π + π/7) 
= (-1/8). ( 1/ sin π/7 ). ( - sin π/7 ) 
= 1/8
 
Regards
Arun (askIITians forum expert)